Two point particles with charges 7.29μC and 5.87μC are held in place by 10.25-N forces on each charge in appropriate directions. Find the distance between the charges.


Round your answer to 3 decimal places


1
Expert's answer
2021-05-09T13:07:42-0400

"F=k\frac{q_1q_2}{r^2}\to r=\sqrt{\frac{kq_1q_2}{F}}=\sqrt{\frac{9\cdot10^9\cdot 7.29\cdot10^{-6}\cdot 5.87\cdot10^{-6}}{10.25}}=0.194\ (m)" . Answer